# REP: Resource-Efficient Prompting for Rehearsal-Free Continual Learning

### v2

@@ 3.2.3 Adaptive Layer Dropping (ALD)
To strengthen our claim, we conduct a simple experiment that drops 25% of layers from various positions. We compare their training wall-clock GPU time and final average accuracy in Figure 4 (See Section 4 for details on these metrics), using L2P with the ViT-L backbone on Split ImageNet-R. Specifically, Top-layer Drop and Bottom-layer Drop statically drop the first and last 25% of layers, respectively, while Random Drop randomly skips 25% of layers across the network. For comparison, we also include No Drop, which performs no layer dropping. Although all these naïve strategies incur accuracy loss, Bottom-layer Drop performs significantly worse than the others. Also, it appears undesirable to bypass only deep layers, even though these layers mainly capture general, task-agnostic features. While Random Drop most closely approaches No Drop’s accuracy, its uniform layer dropping still sacrifices some accuracy. This is why our method ALD carefully accounts for each layer’s spatial aspect in a non-uniform manner to minimize the impact on accuracy.
@@
$\alpha(l)$ reflects the extent of token merging performed by AToM in layer $l$. $n(l)$ represents the original number of tokens, and $n^{\prime}(l)$ is the number of tokens remaining after merging. When the number of merged tokens surpasses the threshold $\tau$, ALD adjusts the layer-dropping probability based on $\alpha$. Since deeper layers tend to merge more tokens with AToM, ALD is more likely to exceed $\tau$ in deeper layers and drop more aggressively. These parameters should be tuned to balance between efficiency and the need to process the nuanced information contained in the merged tokens. We set $\alpha$ to 0.9, and $\tau$ to 16 for ViT-L, 12 for ViT-B, respectively.
One might contemplate exploiting stochastic depth [12], which gradually increases dropping probabilities for deeper layers without a temporal schedule. We compared ALD and stochastic depth and found that ALD generally offers 1.8–3.3% higher accuracy. This may be attributed to early training iterations after task insertion, which play a critical role in stabilizing training losses [21].
Resource Efficiency. REP triggers AToM and ALD for each task insertion to optimize both time and memory usage. Much like our prompt selection optimization, AToM consistently enhances time and memory efficiency. In contrast, ALD solely contributes to reducing time costs because it operates layer-keeping probability $\theta_{t,l}$ starting at 1.0, i.e., no layer dropping, which necessitates traversing full layers.
4 Experiments of REP
Data Generation. To organize task streams, we use three image classification datasets: CIFAR-100 (100 classes) [16], ImageNet-R (200 classes) [11], and PlantDisease (38 classes) [24]. Out of these datasets, ImageNet-R is known for exhibiting much higher intra-class variability among images and an uneven class-size distribution. We divide CIFAR-100 and ImageNet-R into 10 tasks to create Split CIFAR-100 (i.e., 10 classes per task) and Split ImageNet-R (i.e., 20 classes per task), respectively [36, 35, 29]. For PlantDisease, we drop 3 plant disease classes with very few images and organize the remaining 35 classes into 7 tasks to create Split PlantDisease (i.e., 5 classes per task).
Methods. We implement L2P [36], DualPrompt [35], and CODA-Prompt [29] as representative ViT-based prompting methods using PyTorch. L2P and DualPrompt select suitable prompts via inference, similar to [31], while CODA-Prompt employs a strategy for adaptively optimizing prompts without prompt selection, similar to [13]. These methods capitalize on ImageNet pre-trained models as backbones: ViT-L (307M), ViT-B (86M), and ViT-Ti (5.8M)—ViT-Ti is one of the smallest vision transformer models to our knowledge. We categorize the prompt-based methods into three groups based on their memory consumption, as detailed in Table 1, to reflect diverse device memory capacities.
To demonstrate the effectiveness of prompting methods, we also incorporate two state-of-the-art CNN-based methods, BudgetCL [6] and MEMO [40]. Both BudgetCL and MEMO improve model accuracy by leveraging spare memory to store and replay past samples, with MEMO additionally storing model checkpoints from history. These methods also utilize ImageNet pre-trained models as backbones: ResNet-10 (RN10; 5M), ResNet-18 (RN18; 11M), ResNet-26 (RN26; 14M), ResNet-34 (RN34; 22M), and ResNet-101 (RN101; 43M) models. For fair comparisons, we similarly group methods based on their memory usage. We refer readers to supplementary materials for detailed settings of all methods mentioned above.
Hardware and Metrics. Given that not all CL methods are currently feasible on small devices due to memory limitations, we use NVIDIA RTX 3090 Ti as a reference GPU and use the wall-clock time required to complete all tasks—Training wall-clock time correlates linearly with energy usage [21]. In supplementary materials, we evaluate our memory-efficient method directly on NVIDIA Jetson TX2, although it is applicable beyond this specific device. We also report the final average accuracy by averaging the accuracy of all classes after the last task training, and the forgetting metric [5, 29], which assesses the decline in task accuracy averaged over all tasks.
Figure 6: REP over the best methods across various memory budgets and datasets.
4.1 Preliminary Empirical Study
$\bullet$ Performance scaling. ViT-based methods scale well, with larger backbone networks overall yielding higher accuracy. In contrast, CNN-based methods scale poorly with increased backbone size. Specifically, on Split ImageNet-R, a more challenging dataset, MEMO/RN34 is only 2.95% better than MEMO/RN18, despite being around twice larger. On Split CIFAR-100, increasing the backbone size yields only limited gains.
$\bullet$ Memory efficiency. CNN-based methods are often considered more memory-efficient because they use backbones a magnitude smaller than ViT models. Thus, it is commonly believed that ample memory can be allocated to replay samples or past model checkpoints [40], which help improve model performance. However, in CL, which involves typical training iterations, a substantial amount of memory must be allocated for feature maps [20], leaving little room for memory buffers. For instance, for a device memory range of 1–4GB in Table 1, the backbone of MEMO/RN18 consumes only 43MB of memory (6.5% of ViT-L), but feature maps occupy as much as 1.7GB (more than ViT-based methods). This causes MEMO/RN18 to be short on memory for the replay buffer, dramatically degrading model accuracy, as observed in  Figure 5.
4.2 REP on Best-Performing Methods
Next, in each graph in Figure 6, we demonstrate the efficacy of REP when applied to the best-performing CL method for each combination of memory budget and dataset observed in Figure 5. Accuracy is indicated by stars ($\bigstar$), while GPU wall-clock time is indicated by bars ($\blacksquare$). The memory usage in GB is noted on top of the bars. Please refer to the supplementary document for the full results of all datasets.
Resource Efficiency. The best-performing baselines optimized with REP demand significantly smaller amounts of system resources. Specifically, REP reduces memory consumption by 33% for L2P/ViT-L, 11% for DualPrompt/ViT-B, and 5% for DualPrompt/ViT-Ti. Despite the frozen backbone not requiring gradients for parameter updates, it still performs the backward pass to calculate classification and prompt losses for updating learnable parameters. Therefore, all intermediate data are stored in memory, similar to training a non-frozen model. This memory usage is reduced by applying AToM and ALD. In addition, REP exhibits a 16–47% reduction in GPU wall-clock time compared to competing methods, leading to great energy savings.
Accuracy. L2P, DualPrompt, and CODA-Prompt were initially evaluated using only ViT-B backbone, focusing primarily on accuracy in their original studies. In comparison, our evaluation extends these methods across diverse backbones while taking into account their energy-accuracy trade-offs in on-device CL across various memory constraints. When augmented with our techniques, these methods show significant improvements in resource efficiency with minimal accuracy loss across all memory budgets: 0.09–1.90% for Split PlantDisease, 0.03–1.07% for Split CIFAR-100, and 0.26–0.86% for Split ImageNet-R. Especially for the Split PlantDieases, REP with DualPrompt/ViT-B performed best under both 1–4GB and 4–8GB budgets.
4.3 Ablation and Additional Study
@@
REP (ours) 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
(1) $f_{\text{efficient}}$ 74.79$\pm$0.78 4.04$\pm$0.80 3698.78 5.53
(2) AToM 74.91$\pm$0.77 3.47$\pm$0.84 4442.04 5.60
(3) ALD 74.53$\pm$0.78 3.63$\pm$0.67 4243.97 6.67
(4) $f_{\text{efficient}}$ + AToM 74.15$\pm$0.82 4.03$\pm$0.76 2861.35 4.46
(5) $f_{\text{efficient}}$ + ALD 74.52$\pm$0.43 3.43$\pm$0.54 3206.74 5.51
(6) AToM + ALD 74.57$\pm$0.37 2.56$\pm$0.52 3448.76 4.77
Table 2: Ablating REP’s components. They contribute to resource efficiency and accuracy.
@@
REP (ours) 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
(1) w/ ToMe 70.19$\pm$0.74 2.63$\pm$0.92 2917.84 3.74
(2) w/ PLD 73.33$\pm$0.71 3.88$\pm$0.69 2741.90 4.46
2 76.16$\pm$0.74 3.54$\pm$0.68 2808.43 5.48
4 75.32$\pm$0.32 3.33$\pm$0.45 2714.04 5.16
8 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
0.25 73.18$\pm$0.55 4.03$\pm$0.49 2362.07 4.45
0.50 75.34$\pm$0.86 3.58$\pm$0.29 2542.43 4.45
0.75 74.44$\pm$0.60 3.49$\pm$0.53 2861.35 4.45
Component Ablation. In  Table 2, we ablate REP ’s components to assess their contributions to resource efficiency. Each component plays a key role in reducing memory and compute time, with ALD affecting compute time only, as expected. Notably, AToM substantially optimizes both resource usages.
Algorithm Validation. Token merging (ToMe) [3] and progressive layer dropping (PLD) [38] are specifically designed to accelerate traditional transformer-based model training. To validate the importance of incorporating our adaptive techniques instead in CL, we evaluate how REP performs in case AToM and ALD are replaced with ToMe and PLD, respectively. The results are presented in Table 3. Although applying ToMe or PLD improves system efficiency over L2P, it results in an accuracy loss of 5.15% or 2.01%, respectively, compared to using our techniques. This indicates that ToMe and PLD are less desirable for achieving our goal, i.e., enhancing system efficiency without compromising accuracy.
AToM and ALD Intensity. Table 4 shows the effects of varying intensities of AToM and ALD, focusing on the number of merged tokens ($n$) in AToM and the keep ratio ($\theta$) in ALD. AToM appears to maintain stable accuracy even as more tokens are merged. In contrast, for ALD, a lower keep ratio improves system efficiency by reducing iteration time and memory usage, but it can markedly impair model accuracy if the ratio is too low. Overall, when used with caution, AToM and ALD can effectively balance system efficiency and accuracy.
Longer Task Sequence. We examine how REP performs on a longer task sequence of 20 tasks. A longer task sequence represents a more challenging scenario, where performance drop as the number of tasks increases can become more pronounced. We observe that REP consistently preserves model accuracy (from 72.77$\pm$0.19 in L2P to 72.58$\pm$0.59 in REP) while improving resource efficiency by reducing GPU wall-clock time and memory usage by 47.3% and 33%, respectively.

### v3

@@ Appendix A Algorithm Details
Appendix B Implementation Details
Unless otherwise stated, we use the term REP-L to denote REP applied to L2P with a ViT-L backbone, which stands out as the best-performing baseline method for Split CIFAR-100 and Split ImageNet-R within the 4–8GB memory range. Similarly, the notations L, B, and Ti are used to generally denote ViT-L, ViT-B, and ViT-Ti backbones, respectively, in the prompt update stage of any prompt-based method.
B.1 Prompt-Based Methods for REP
We present details on the three major prompt-based methods to which REP is integrated for resource efficiency.
Learning to Prompt (L2P) positions prompts at the first layer of the transformer architecture [39]. These prompts are learnable parameters that dynamically evolve with the training process. The mechanism begins with a prompt pool $P=\{p_{1},p_{2},...,p_{m}\}\subset\mathbb{R}^{L_{p}\times D}$.
For a given input $x_{i}^{j}$ in task $T_{i}$, L2P computes a query feature $q(x_{i}^{j})\in\mathbb{R}^{D}$ to select the corresponding prompt. The prompt $p^{*}$ is selected based on maximizing the cosine similarity with respect to the query:
$p^{*}=\underset{p_{k}\in P}{\mathrm{argmax}}\sum_{c=1}^{L_{p}}\frac{\big{\langle}q(x_{i}^{j}),[p_{k}]_{(c,:)}^{\top}\big{\rangle}}{\|q(x_{i}^{j})\|\|[p_{k}]_{(c,:)}^{\top}\|},$ (7)
where $[p_{k}]_{(c,:)}$ is the $c$-th row of $p_{k}$.
The selected prompt $p^{*}$ is concatenated with the input embedding $z(x_{i}^{j})$ to form the prompt-augmented input $z^{\prime}(x_{i}^{j})=[p^{*};z(x_{i}^{j})]$. The training objective of L2P balances classification loss $L_{\text{class}}$ and prompt-adjustment loss $L_{\text{prompt}}$:
$L=L_{\text{class}}(z^{\prime}(x_{i}^{j}),y_{i}^{j})+L_{\text{prompt}}(p^{*},q(x_{i}^{j})).$ (8)
DualPrompt leverages prompts at multiple layers of the transformer architecture [38]. It introduces a general prompt $g\in\mathbb{R}^{L_{g}\times D}$ and a set of task-specific prompts $E=\{e_{1},e_{2},...,e_{T}\}\subset\mathbb{R}^{L_{e}\times D}$. These prompts are incorporated at specified layers in the transformer model.
For a given input $x_{i}^{j}$ in task $T_{i}$, the model’s transformer layers $f$ are modified by attaching $g$ and $e_{i}$ to the layers, resulting in a prompted architecture $f_{g,e_{i}}$. The feature transformation $h_{i}^{j}$ for the input sample $x_{i}^{j}$ is then obtained as:
$h_{i}^{j}=f_{g,e_{i}}(x_{i}^{j}).$ (9)
Similar to L2P, DualPrompt optimizes $L_{\text{class}}$ and $L_{\text{prompt}}$:
$L=L_{\text{class}}(h_{i}^{j},y_{i}^{j})+L_{\text{prompt}}(g,e_{i}).$ (10)
CODA-Prompt decomposes learnable prompts into components and uses an attention mechanism from a pre-trained ViT model to select relevant prompts [31]. Instead of a single prompt, CODA-Prompt learns a set of prompt components $P=\{P_{1},P_{2},...,P_{M}\}$. The final prompt $p$ is a weighted sum:
$p=\sum_{m}\alpha_{m}P_{m},$ (11)
where weights $\alpha$ are determined based on the query $q(x)$ and keys $K\in\mathbb{R}^{D\times M}$:
$\alpha=\gamma(q(x),K).$ (12)
When the task changes, the current components are frozen, and new one is added, promoting orthogonality:
$L_{\text{ortho}}(B)=\|BB^{\top}-I\|^{2},$ (13)
where $B$ represents $P$, $K$, or $A$ ($A$ represents the attention vector). The full optimization target is:
$\min_{P^{n},K^{n},A^{n},\phi^{n}}$ $\,L\left(f_{\phi}(f_{\theta,P,K,A}(x)),y\right)+\lambda\left(L_{\text{ortho}}(P)+L_{\text{ortho}}(K)+L_{\text{ortho}}(A)\right)$ (14)
where $P^{n}$, $K^{n}$, and $A^{n}$ are new components, and $\lambda$ balances the orthogonality loss [31].
B.2 Hyperparameters and Configurations
For all prompt-based CL methods, we employ the ADAM optimizer [17] configured with hyperparameters $\beta_{1}=0.99$ and $\beta_{2}=0.999$. In contrast, BudgetCL and MEMO adopt the stochastic gradient descent (SGD) optimizer with a momentum coefficient of 0.9 and a weight decay of 0.002, both validated as effective by similar replay buffer-based methods [39]. In REP, we set the prompt pool size to 10 and the prompt length to 5, following the original implementations of L2P [^2] L2P PyTorch Implementation. https://github.com/JH-LEE-KR/l2p-pytorch. and DualPrompt. [^3] DualPrompt PyTorch Implementation. https://github.com/JH-LEE-KR/dualprompt-pytorch. The learning rate for the prompt-based methods is set to 0.001875. For CODA-Prompt, we implement a cosine decay learning rate strategy, as outlined in its original paper [31]. For BudgetCL and MEMO, the learning rate is set to 0.1 and decays by 0.1 at 80 and 150 epochs following the original paper [43].
We consistently use the mini-batch size of 16 for all methods to maintain a uniform computational load for each training iteration. For fair comparisons, each method performs 1875, 1080, and 2583 iterations per task insertion for Split CIFAR-100 [18], Split ImageNet-R [12], and Split PlantDisease [26], respectively. This setup ensures that among prompt-based methods, iteration time itself correlates linearly with energy usage, as training wall-clock time similarly reflects energy consumption as discussed in the main paper.
BudgetCL and MEMO select 8 samples from the new task and another 8 samples from old tasks in the replay buffer to organize 16 samples in each mini-batch. We adopt the setup in [43] and train 200 epochs for all datasets used. Thus, BudgetCL and MEMO perform more iterations per epoch for a bigger replay buffer, translating into longer GPU time and more energy usage.
Appendix C Competing Methods with Original Setups
Table 6: Results on Split CIFAR-100, with ViT-B as the backbone model (following the original setups).
Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP-L (ours) 86.89$\pm$0.57 4.65$\pm$0.14 4420.99 4.45
L2P-B 84.36$\pm$0.68 7.35$\pm$0.38 2802.42 14.20
DualPrompt-B 85.27$\pm$0.75 5.16$\pm$0.09 2624.06 12.30
CODA-Prompt-B 86.09$\pm$0.73 1.67$\pm$0.26 3226.58 18.83
Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP-L (ours) 75.34$\pm$0.16 3.58$\pm$0.29 2542.43 4.45
L2P-B 59.66$\pm$0.81 9.73$\pm$0.47 1609.88 14.20
DualPrompt-B 67.45$\pm$0.88 4.68$\pm$0.20 1507.42 12.30
CODA-Prompt-B 75.52$\pm$0.74 1.64$\pm$0.10 1853.54 18.83
For the evaluation of L2P, DualPrompt, and CODA-Prompt in the main paper, we adhere to the hyperparameters specified in the original studies, except for the number of iterations per task and batch size. We adopt a smaller batch size to account for on-device memory limitations, which prevent the use of the original batch sizes. Nonetheless, we present the results of these three methods using their original hyperparameters in Table 6 and Table 7, respectively. Although these methods are based on the ViT-B backbone, they consume significantly more memory than REP-L (with the ViT-L backbone), primarily due to their originally large batch sizes.
Appendix D Additional Datasets
Table 8: Results on Split CUB-200, with ViT-L as the backbone model.
Method Acc. ($\uparrow$) Fgt. ($\downarrow$) GPU Time (s) Mem. (GB)
REP-L (ours) 74.70$\pm$1.04 6.67$\pm$0.34 529.67 4.45
L2P-L 74.70$\pm$1.02 6.68$\pm$1.35 1005.57 6.67
DualPrompt-L 72.35$\pm$0.98 7.82$\pm$0.31 953.42 5.93
CODA-Prompt-L 79.46$\pm$0.84 5.84$\pm$0.99 1257.82 13.95
We also expand our experiments on REP by incorporating an additional dataset: Split CUB-200 [33], which is designed for fine-grained image classification. Split CUB-200 contains 5,994 bird images categorized into 200 classes, which are split into 5 CL tasks, each having 40 classes. We present the experimental results for all prompt-based methods using ViT-L as the backbone in Table 8. REP-L outperforms competing methods in resource efficiency by bringing significant reductions in both computation time and memory usage. Importantly, REP-L is on par with its baseline model L2P in accuracy. Note that while adopting CODA-Prompt to implement REP can deliver better accuracy, its memory usage far exceeds all memory ranges under consideration.
Appendix E Extended Ablation and Additional Study
As the paper mainly discusses the ablation study based on Split ImageNet-R, we here present an extended ablation study of REP using Split CIFAR-100. The results are shown in Table 9. Similar to the findings from Split ImageNet-R, the ablation of any component within REP for the Split CIFAR-100 dataset impacts resource efficiency. All components contribute to the reduction of computation time and memory usage, with AToM particularly standing out in both aspects of resource efficiency.
Furthermore, we perform algorithm validation on Split CIFAR-100, for which we present the results in Table 10. Consistent with the Split ImageNet-R results, using static token merging (ToMe) [3] or progressive layer dropping (PLD) [41] instead of adaptive token merging (AToM) or adaptive layer dropping (ALD) negatively affects model accuracy on the Split CIFAR-100 dataset. This experiment further corroborates the efficacy of our proposed algorithms in optimizing resource usage without compromising model accuracy in CL scenarios.
Table 9: Extended component ablation for Split CIFAR-100. Ablating any component of REP results in lower resource efficiency by increasing training time and memory consumption.
@@ Appendix E Extended Ablation and Additional Study
REP-L (ours) 86.89$\pm$0.57 4.65$\pm$0.14 4420.99 4.45
(1) $f_{\text{efficient}}$ 86.38$\pm$0.43 5.82$\pm$0.81 6096.14 5.53
(2) AToM 86.43$\pm$0.23 4.94$\pm$0.77 7724.21 5.60
(3) ALD 86.79$\pm$0.73 5.01$\pm$0.39 7379.79 6.67
(4) $f_{\text{efficient}}$ + AToM 85.97$\pm$0.54 4.99$\pm$0.73 4975.57 4.46
(5) $f_{\text{efficient}}$ + ALD 86.07$\pm$0.18 4.70$\pm$0.66 5576.16 5.51
(6) AToM + ALD 86.08$\pm$0.48 3.97$\pm$0.42 5997.01 4.77
Table 10: Extended algorithm validation for Split CIFAR-100. Changing our adaptive algorithms to conventional static token merging (ToMe) or progressive layer dropping (PLD) harms model accuracy.
@@
REP-L (ours) 86.89$\pm$0.57 4.65$\pm$0.14 4420.99 4.45
(1) w/ ToMe 83.55$\pm$0.14 3.73$\pm$0.61 5073.80 3.74
(2) w/ PLD 84.40$\pm$0.35 4.88$\pm$0.83 4767.86 4.46
Appendix F Extended Preliminary Empirical Study
Figure 6: Extended analysis of the energy-accuracy trade-off on Split PlantDisease dataset. While CNN-based methods may excel on certain memory ranges, ViT-based methods consistently demonstrate better resource efficiency.
We present an extended analysis of the energy-accuracy trade-off using the Split PlantDisease dataset [26] in Figure 6. Interestingly, CNN-based methods can achieve higher accuracy on this dataset, albeit with higher computational costs. Since the PlantDisease dataset consists of localized leaf images with various disease patterns, it appears to favor CNN architectures like ResNets, which are excellent at capturing local features. On the contrary, ViT-based methods, which excel at capturing global information due to their pre-trained weights, may not perform very well on tasks requiring extensive localized feature extraction. Nonetheless, ViT-based methods demonstrate significantly greater resource efficiency. Thus, for this specific benchmark, the choice between CNN-based methods and ViT-base methods simply depends on the priority: higher resource efficiency vs. higher accuracy.
Appendix G REP under Varying Compute Budgets
Table 11: Impact of REP under various computation budgets (i.e., number of iterations per task).
Method 313 625 938 1250 1563 1875
REP-L (ours) 84.06 85.47 86.60 86.66 86.57 86.89
L2P-L 84.36 86.28 87.34 87.76 87.83 88.24
We compare REP-L with L2P across various compute budgets defined by the number of iterations per task. The comparison results using Split CIFAR-100 are shown in Table 11. Evidently, the accuracy of REP-L is comparable to that of L2P across six compute budgets, with the performance gap getting increasingly marginal at lower compute budgets. We observe that the overall accuracy for both methods does not drop significantly with reduced compute budgets. This is attributed to the fact that Split CIFAR-100 is considered a less complex CL benchmark.
Appendix H REP on Edge Devices
Miro [23] introduces a dynamic approach for fine-tuning key design parameters of rehearsal-based CL methods to achieve high model accuracy while simultaneously reducing energy costs, i.e., high cost-effectiveness. Miro’s methodology centers on identifying the optimal memory size within the device’s capacity to accommodate both old and new samples for training. We compare REP with Miro directly on a reference edge device NVIDIA Jetson TX2444NVIDIA. Jetson TX2 Module. https://developer.nvidia.com/embedded/
jetson-tx2.
. This device is equipped with a 256-core NVIDIA Pascal GPU, 8GB of RAM, and a 32GB eMMC 5.1 drive. The RAM is a single unified memory shared by the CPU and GPU. We report energy usage as joules (J) obtained by multiplying power by time. Power usage is measured by reading built-in sensor values provided by Jetson devices for the GPU, RAM, CPU, and I/O connection [^5] Convenient Power Measurements on the Jetson TX2/Tegra X2 Board. https://embeddeddl.wordpress.com/2018/04/25/convenient-power-measurements-on-the-jetson-tx2-tegra-x2-board/..
H.1 Power Usage Breakdown
Figure 7: Power breakdown for training a ViT-B model on NVIDIA Jetson TX2.
When implementing CL on edge devices, the majority of power consumption that influences energy usage during the training of a new task comes from GPU operations. Figure 7 shows power consumption on NVIDIA Jetson TX2 across individual system components. Static power is measured when the system is inactive, while dynamic power is measured during the training of a ViT-B model. During training, power usage surges up to 6.5$\times$, with the GPU contributing 60% to the dynamic power.
H.2 REP vs Miro
In this experiment, we explore energy-accuracy trade-offs of REP-L and Miro. Figure 8 visually represents the comparison results. In each graph, the x-axis signifies total energy usage (lower is better), while the y-axis signifies final average accuracy (higher is better). Hence, a more cost-effective strategy is positioned closer to the upper-left corner of the graph. Regarding Miro, we maintain the hyperparameters as suggested in the original work [23] but incorporate the use of pre-trained ResNet-18 (RN18; 11M), ResNet-34 (RN34; 22M), and ResNet-50 (RN50; 25M)[11] models instead of non-pre-trained ones to enhance overall performance. For REP-L, we vary the number of training iterations per task insertion to match the energy usage of Miro variants with different ResNet models.
Table 12: Comparison of REP-L with ADAM variations.
Method Acc. ($\uparrow$) GPU Time (s) Mem. (GB)
REP-L (ours) 72.58 2542.43 4.45
VPT-Deep-L 66.70 1444.62 12.61
VPT-Shallow-L 64.53 1271.69 6.89
SSF-L 72.16 1573.70 22.09
Adapter-L 62.05 680.05 12.01
Table 13: Results on Split ImageNet-R, with ViT-L as the backbone model.